Determine the [H+] and [OH-] in
(a) 5.0 x 10-5 M HNO3
(b) 2.0 x 10-6 M KOH
1
Expert's answer
2020-05-24T03:00:59-0400

(a) As HNO3 is a strong acid, it dissociates completely. As a result, [H+] = [HNO3] = 5.0 × 10-5 M. As [H+] × [OH-] = 1.0 × 10-14 :

[OH-] = 1.0 × 10-14 / [H+] = 1.0 × 10-14 M / 5.0 × 10-5 = 2.0 × 10-10 M

(b) As KOH is a strong acid, it dissociates completely. As a result, [OH-] = [KOH] = 2.0 × 10-6 M.

[H+] = 1.0 × 10-14 / [OH-] = 1.0 × 10-14 / 2.0 × 10-6 = 5.0 × 10-9 M


Answer: (a) [H+] = 5.0 × 10-5 M, [OH-] = 2.0 × 10-10 M; (b) [OH-] = 2.0 × 10-6 M, [H+] = 5.0 × 10-9.